OpenStudy (anonymous):

What value should be added to the expression to create a perfect square? x power of 2 - 20x -25 25 100 -100

OpenStudy (anonymous):

100 x^2 - 20x + 100 = (x - 10)^2 ax^2 + 2abx + b^2 = (ax + b)^2

OpenStudy (maheshmeghwal9):